OpenStudy (anonymous):

A right circular cylinder is inscribed in a cone of height h and radius r. Determine the dimension of the cylinder with the largest possible volume.

OpenStudy (anonymous):

Draw a diagram of the cross-section. If the height and radius of the cylinder are h and r respectively, then from similar triangles (or taking tan of the semi-vertical angle r/(H-h)=R/H so h=H(1 - r/R) The vol of the cylinder V=pir^2h=pir^2H(1-r/R)=piH(r^2 -r^3/R) This is max when r^2-r^3/R is max which is when 2r-3r^2/R =0 (dV/dr=0) giving r=2R/3 and h=H/3

OpenStudy (anonymous):

So would the cross section of a cylinder be a rectangle?

OpenStudy (anonymous):

i need go sorry

OpenStudy (anonymous):

ok thanks anyways